OpenStudy (anonymous):

An electron inside an electric field region between two parallel plates as shown below is rejected upward at a speed of Vo=5.83x10^6 m/s and at an angle of teta=39.0 degrees ; E=1870 N/C (direct upward), d =1.97cm and L=6.20cm. Will the electron strike one of the two plates? If yes, which plate will it strike and what distance from the left edge?

OpenStudy (anonymous):



OpenStudy (anonymous):

It's parabolic motion, you can use x=v0*cos(39.0)*t and y=vo*sin(39.0)*t-1.2 *(eE/me) *t^2 vy=v0*sin(39.0)-(eE/me) *t And here's strategy to find the answers:

OpenStudy (anonymous):

but no strategy to find e strike the plat